# A Hierarchical Approach to Quantum Many-Body Systems in Structured Environments

### v1

@@ Abstract
Cavity quantum materials combine the rich many-body physics of condensed matter systems with strong coupling to the surrounding electromagnetic field, which presents both novel prospects and intricate challenges. One is often interested in the properties of one specific aspect of the material, e.g. the electronic many-body dynamics, subject to a structured bath of phononic and photonic modes. Open quantum systems featuring non-Markovian dynamics are routinely solved using techniques such as the Hierarchical Equations of Motion (HEOM) but their usage of the system density-matrix renders them intractable for many-body systems. Here, we combine the HEOM with the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy to reach a consistent and rigorous description of open many-body systems and their quantum dynamics. We demonstrate first the strength and limitations of this stacked hierarchy for superradiant emission and spin-squeezing of established quantum optical models before presenting its full potential for quantum many-body systems. In particular, we explicitly simulate the impact of charge noise on the dynamic of the Fermi-Hubbard model subject to a structured bath comprising cavity and vibro-phononic environment. Strong optical coupling not only modifies the dynamic of the many-body system but serves furthermore as measurement channel providing information about the correlated motion imprinted by charge noise. Our work establishes an accessible, yet rigorous, route between condensed matter and quantum optics, fostering the growth of a new domain at their interface.
Please note, this early version will be refined and updated with additional results in the near future.
@@ I Introduction
The ability to control many-body quantum systems with quantized light is imperative in modern quantum technologies such as photonics, quantum memories, quantum sensing, or quantum information.[1, 2, 3, 4] An idealized description of the fundamental building block, the quantum emitter, is challenged by disturbances, such as charge noise or phonon scattering, that hint at the inherent many-body character of each quantum system. Strong light-matter coupling allows to control interactions over extended length scales[5, 6, 7, 8, 9], holding promise to design materials [10, 11, 12], devices [13, 14], or even chemical properties [15, 16, 17, 18, 19] non-intrinsically and on demand. The many-body physics originating from electronic and vibrational structure, as well as their respective decoherence, is crucial for a comprehensive understanding and refinement of this control-strategy.[20, 21, 22, 23, 24]
@@
Our primary challenge in this study is to address physical processes wherein many-body interactions, strong light-matter coupling, and strong system-environment coupling are all fundamental. Achieving this goal necessitates combining concepts from quantum optics and solid-state physics. Quantum optics often treats light-matter interaction non-perturbatively but simplifies the structure of materials drastically. Conversely, condensed matter or quantum chemistry approaches accurately describe the material system but oversimplify interactions with surrounding environments. Our aim is to bridge quantum optical and condensed matter/quantum chemistry approaches to create a framework for studying the effects of strong light-matter coupling on many-body physics more efficiently than previously possible.
One approach to handle non-Markovian dynamics is to resort exact numerical methods, where the effects of the environment are taken into account stochastically. These methods could be wave function based, such as non-Markovian Quantum State Diffusion (NMQSD) [25, 26, 27] or density matrix based such as the Hierarchical Equations Of Motion (HEOM) [28, 29]. In this work we focus on the latter approach. HEOM has been used to solve a large variety of problems such as the dynamics of the Holstein polaron [30] or the problem of coupling to multiple environments [31]. To use HEOM, the bath correlation function, which measures the response of the environment to the open system needs to be fitted with exponential functions [32]. The resulting hierarchy couples the physical density matrix to auxiliary density matrices, and the hierarchy’s depth depends on the system-environment coupling strength. However, this approach becomes impractical for many-body systems due to the exponential scaling of physical and auxiliary states with system size. Our aim is to combine the HEOM approach with an efficient strategy for many-body dynamics.
The second challenge to face is the strong coupling between the material system and the cavity field. Micro- and plasmonic cavities will often require the consideration of multiple modes [33, 13]. Commonly used approaches, such as adiabatic elimination [34] or mean field [35], show at best limited success at approximating the exact dynamic. Cluster expansion techniques for the light matter interaction [36, 37] and HEOM related expansions for the material system alone have been proposed [38, 39] but they still suffer from poor scaling for many-body quantum systems. Entering the ultrastrong coupling regime [40], counter-rotating [41] and self-polarization terms [42, 43] become essential which necessitates the usage of the correct Lindblad dissipator that connect between the correlated eigenstates. In this work, we will limit ourselves to the strong coupling regime where such effects play a subordinate role.
@@
Lastly, the third challenge is to control the exponential scaling of the many body system. Wavefunction-based approaches, such as tensor networks [44] or matrix product states, increase exponentially in cost over time which limits their use to short time-scales [45, 46]. Non-equilibrium Green functions approaches made recent progress after the successful integration of the generalized Kadanoff-Baym ansatz [47, 48] but a consistent treatment of complex structured baths remains challenging. An attractive approach with good scaling properties is to perform a cluster expansion of the density matrix into single-,two-,three- and $N$-body reduced density matrices and then truncate the resulting Bogoliubov–Born–Green–Kirkwood–Yvo (BBGKY) hierarchy at some low order [49, 50].
@@
Here, we combine BBGKY and HEOM to establish a stacked BBGKY-HEOM hierarchy that provides an efficient framework for the simulation of correlated many-body systems subject to structured baths. The remainder of this article is structured as follows: Sec. II provides a short introduction to non-Markovian open-system and many-body theory before the derivation of the BBGKY-HEOM equations in Sec. II.2. We illustrate their performance, limitation, and associated physical intuition for various systems in Sec. II. In particular, Sec. III.1 demonstrates that BBGKY-HEOM succeeds in reproducing the quantum many-emitter physics in the superradiant driven Tavis-Cummings model while retaining constant scaling with emitter size. We subsequently shift our focus to many-body electronic systems in the form of the Fermi-Hubbard model coupled to a lossy cavity mode and subject to charge noise in Sec. III.2. This system in then embedded into a highly structured bath in Sec. III.3 that represents phononic motion in organic crystals. The BBGKY-HEOM provides accurate predictions for as long as the onsite repulsion $U$, and with it the degree of electronic correlation, remains moderate. We notice that the bath adds a stabilizing contribution to the BBGKY equations. Sec. IV finally concludes our study and provides an outlook into the nearer future.
@@ II Theory
A material comprising electrons and nuclei moving at non-relativistic velocities is described by Schrödinger’s equation. Its interplay with electromagnetic fields necessitates the consideration of the (quantized) normal modes of the Maxwell equations. This combined system can be described using the minimal coupling Hamiltonian. In Coulomb gauge,
@@
for $N_{e}$ electrons and $N_{n}$ nuclei with charge $q_{i}$ and mass $m_{i}$. We will use atomic units in the following. Eq. (1) treats the transversal degrees separate from the longitudinal and instantaneous Coulombic interactions $H_{\parallel}$. The latter includes electronic and nuclear interactions which are typically treated in first quantization at first
@@
Their interplay results in rich many-body physics that gives rise to electronic, vibronic, and phononic structure and, unfortunately, to an exponentially increasing complexity. The system-bath strategy aims now at identifying the smallest possible subsystem that we can treat explicitly while absorbing the remainder into a bath[51]. First step is to simplify the Hamiltonian to a form that contains only linear interactions between various subsystems
@@
Such a separation can be formalized using the projection operators $P$ and $Q=1-P$, which are defined in such a way that the relevant state of the subsystem is obtained by projecting the total state $\rho_{tot}(t)$ to the relevant subspace
$\rho_{sys}(t)=P\rho_{tot}(t).$ (4)
[…]
The gauge freedom of the electromagnetic modes provides various different paths. An intuitive approach is to isolate the bilinear term $\sum_{i}-i\hbar q_{i}/m_{i}\hat{\textbf{A}}(\textbf{r}_{i})\cdot\boldsymbol{\nabla}_{i}$ as coupling operator and absorb the quadratic diamagnetic term into dressed photonic operators[42, 43]. A second direction, often more promising for finite systems, is to shift into the multi-polar gauge (detailed in App. C). Combination with the long-wavelength approximation provides then a simple bilinear coupling between dipole moment and displacement field. In addition, self-polarization terms modify the dynamic of the system and the redefinition of the optical creation and annihilation operators introduce the need to reconsider the interpretation of system and bath. Vibrational and phononic interaction follow similarly. The influence of nuclei motion on the electron-nuclear interaction can be expanded around the equilibrium configuration
$\hat{V}_{en}\approx\hat{V}_{en}|_{0}+\sum_{j}\delta\textbf{R}_{j}\cdot\boldsymbol{\nabla}_{j}\hat{V}_{en}|_{0}+...$ (5)
and truncated at harmonic order, resulting in a bilinearly coupled electron-phonon or electron-vibration coupling that depends on the electronic configuration. Note that also here second-order (known as Debye-Waller) corrections appear that require to carefully scrutinize the separation between system and bath.[52] The following derivations and demonstrations are agnostic to such subtleties as we will focus on the development and illustration of the BBGKY-HEOM methodology. A thorough derivation of system-bath couplings for a specific physical realization is evidently essential for meaningful predictions but it is not the focus of this manuscript.
@@ II.1 Many-body Open Quantum Systems beyond the Markovian Limit
Very often the environment can be modeled as a collection of non-interacting quantum harmonic oscillators. This is for example frequently the case in quantum optics [53], or when the system is coupled weakly to a large number of degrees of freedom where central limit type of arguments yield an approximately Gaussian response [54]. When a description of the environment in terms of harmonic oscillators is suitable, then the HEOM approach yields an exact description for the evolution of the open quantum system [28, 29, 55]. In fact, HEOM type of techniques can be even extended to cases where the environment is anharmonic [56, 57]. In this work, we focus on the case of harmonic baths, either corresponding to optical modes confined in an imperfect cavity or a phonon bath into which a quantum emitter is embedded. The influence functional of the bath can then be computed exactly under the harmonic assumption when using the path integral formalism [58]. HEOM follows from this influence functional [29] and we provide a sketch of the derivation for the open system in App. D.
@@
Following the previous motivation, we start with a many-body system that is linearly coupled to a collection of harmonic oscillators. Especially in the case of cavity quantum materials many different environments (e.g. phonon- and electromagnetic environment) can be relevant for the problem. We label these distinct environments with $k$, such that the bosonic annihilation (creation) operators belonging to mode $\lambda$ of bath $k$ are given by $a_{\lambda}^{k}$ ($a_{\lambda}^{k\dagger}$). The evolution of the total state $\rho_{tot}$ of system and environments is then described by
$\dot{\rho}_{tot}=$ $-i[H,\rho_{tot}],$ (6)
$H=$ $H_{sys}+\sum_{k,\lambda}g_{\lambda}^{k}(L^{k}a_{\lambda}^{k,\dagger}+L^{k\dagger}a^{k}_{\lambda})+\sum_{\lambda}\omega^{k}_{\lambda}a_{\lambda}^{k\dagger}a^{k}_{\lambda},$ (7)
where $L^{k}$ is the system operator that describes the coupling to environment $k$ and $\omega_{\lambda}^{k}$, $g_{\lambda}^{k}$ are the frequencies and coupling strengths of the respective modes. The environments are characterized by their spectral density $J_{k}(\omega)=\sum_{\lambda}\lvert g^{k}_{\lambda}\rvert^{2}\delta(\omega-\omega^{k}_{\lambda})$ or equivalently by their bath correlation function $\alpha_{k}(\tau)=\int_{0}^{\infty}J_{k}(\omega)e^{-i\omega\tau}\mathrm{d}{\tau}$. We are foremost interested in the reduced dynamics in the system Hilbertspace $\mathcal{H}_{sys}$ but it should be noticed that HEOM keeps track of the bath degrees of freedom. For the moment, the $N$ particles of the many-body Hamiltonian are assumed to behave identical under a single particle Hamiltonian $H_{i}$ and experience two-particle interactions $V_{ij}$. In addition, we assume identical coupling to the environments via the single particle operator $L^{k}_{i}$, such that the many-body Hamiltonian and the coupling operators can be decomposed into
$\begin{split}H_{sys}=&\sum_{i=1}^{N}H_{i}+\sum_{i\neq j}^{N}V_{ij},\quad L^{k}=\sum_{i=1}^{N}L^{k}_{i}.\end{split}$ (8)
@@
The evolution equation (6) is therefore invariant under particle exchange. Such assumptions are natural for indistinguishable particles but one can easily relax this condition if one would like to describe for example ensembles of inhomogeneously broadened emitters.
@@ II.2 Combining HEOM with BBGKY
Despite the restrictions employed above, the resulting system-bath Hamiltonian applies to a vast range of systems, from typical quantum optical setups over Rydberg ensembles to molecules in cavities. However, due to the large size of the Hilbert space it is typically not possible to diagonalize Eq. (6) exactly.
Starting point for the derivation of a holistic description that combines many-body and system-bath dynamic is the HEOM equation. We present in the following only the equations for a single environment for the sake of simplicity but note that the extension to multiple environments is trivial and will be illustrated in Sec. III.3. The HEOM equation accounts for the non-Markovianity of the coupled environment with an infinite set of coupled equations for the system density matrices $\rho_{sys}^{(\boldsymbol{n},\boldsymbol{m})}\in\mathcal{H}_{sys}$. Here, $\rho_{sys}^{(\boldsymbol{0},\boldsymbol{0})}=\operatorname{Tr}_{B}(\rho_{tot})=:\rho$ is the reduced density matrix of interest and $\boldsymbol{m},\boldsymbol{n}$ are vector indices labelling (unphysical) auxiliary density matrices. The hierarchy is truncated at a sufficiently high order which is determined by the non-Markovianity of the system-bath interaction, i.e., a Markovian bath corresponds to a truncation at first order according to $\rho_{sys}^{(\boldsymbol{e}_{k},\boldsymbol{0})}\propto L\rho_{sys}^{(\boldsymbol{0},\boldsymbol{0})}$, $\rho_{sys}^{(\boldsymbol{0},\boldsymbol{e}_{k})}\propto\rho_{sys}^{(\boldsymbol{0},\boldsymbol{0})}L^{\dagger}$. The evolution equation for the auxiliary density matrices (including the physical reduced density matrix $\rho_{sys}^{(\boldsymbol{0},\boldsymbol{0})}$) reads
$\begin{split}\dot{\rho}_{sys}^{(\boldsymbol{n},\boldsymbol{m})}&=-i[H_{sys},\rho_{sys}^{(\boldsymbol{n},\boldsymbol{m})}]-\left(\boldsymbol{w}\cdot\boldsymbol{n}+\boldsymbol{w}^{*}\cdot\boldsymbol{m}\right)\rho_{sys}^{(\boldsymbol{n},\boldsymbol{m})}\\ &+\sum_{k=1}^{M}\bigg{(}G_{k}\left(n_{k}L\rho_{sys}^{(\boldsymbol{n}-\boldsymbol{e}_{k},\boldsymbol{m})}+m_{k}\rho_{sys}^{(\boldsymbol{n},\boldsymbol{m}-\boldsymbol{e}_{k})}L^{\dagger}\right)\\ &+\left[\rho_{sys}^{(\boldsymbol{n}+\boldsymbol{e}_{k},\boldsymbol{m})},L^{\dagger}\right]+\left[L,\rho_{sys}^{(\boldsymbol{n},\boldsymbol{m}+\boldsymbol{e}_{k})}\right]\bigg{)},\end{split}$ (9)
where bold symbols indicate the vector character of a variable and $\boldsymbol{e}_{k}$ denotes the unit vector in direction $k$. The complex coefficients $G_{i},w_{i}$ can be obtained from an exponential fit of the bath correlation function $\alpha(\tau)\approx\sum_{i}G_{i}e^{-w_{i}\tau}$ or inferred from scattering theory [59]. The number of coefficients required for the fit sets the dimension of the vector indices and therefore increases the computational complexity.
This equation has been successfully applied to a wide range of systems [55, 60, 61, 62, 63, 64, 65], but a major limitation is the requirement for the entire system density matrix. The latter scales exponentially with the size of the subsystem and is therefore practically inaccessible for the vast majority of many-body quantum systems. Application to many-body systems requires therefore to identify a path that retains the dimensionality of the systems descriptor, thus far $\rho_{sys}^{(\boldsymbol{n},\boldsymbol{m})}$, on a manageable level.
@@
$\begin{split}F_{123}^{(\boldsymbol{m},\boldsymbol{n})}=&N(N-1)(N-2)\operatorname{Tr}_{4,5,...,N}(\rho^{(\boldsymbol{m},\boldsymbol{n})}),\\ F_{12}^{(\boldsymbol{m},\boldsymbol{n})}=&N(N-1)\operatorname{Tr}_{3,4,...,N}(\rho^{(\boldsymbol{m},\boldsymbol{n})}),\\ F_{1}^{(\boldsymbol{m},\boldsymbol{n})}=&N\operatorname{Tr}_{2,3,...,N}(\rho^{(\boldsymbol{m},\boldsymbol{n})}).\end{split}$ (10)
@@
Here $\operatorname{Tr}_{i,i+1,...,N}$ denotes the trace over all particles with index larger than $i$ and the reduced density matrices are normalized to the provide the particle number under contraction [66]. Note that since we are restricting ourselves to permutationally symmetric systems the order of indexing is irrelevant.

### v2

@@ Abstract
Open quantum systems that feature non-Markovian dynamics are routinely solved using techniques such as the Hierarchical Equations of Motion (HEOM). However, their usage of the entire system density-matrix renders them intractable for many-body systems. Here, we combine the HEOM with the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy to achieve a rigorous description of open many-body systems in contact with structured photonic and phononic baths. We first rationalize that this stacked hierarchy accounts for spin-squeezing and superradiant emission despite its applicability to arbitrarily many emitters. The full potential of BBGKY-HEOM is then illustrated for two relevant applications: (i) the explicit treatment of vibrational modes provides access to resonant enhancements in few-emitter lasing, and (ii) the impact of phononic coupling and charge noise on many-body electronic systems embedded in host materials (e.g. molecules in organic crystals) is as relevant as electronic correlation. Our work establishes an accessible, yet rigorous, route between condensed matter and quantum optics, fostering the growth of a new domain at their interface.
@@ I Introduction
The ability to control many-body quantum systems with quantized light is imperative in modern quantum technologies, including photonics, quantum memories, quantum sensing, or quantum information.[1, 2, 3, 4] An idealized description of the fundamental building block, the quantum emitter, is challenged by disturbances, such as charge noise or phonon scattering, that hint at the inherent many-body character of each quantum system. Strong light-matter coupling allows to control interactions over extended length scales[5, 6, 7, 8, 9], holding promise to design materials [10, 11, 12, 13, 14, 15], devices [16, 17, 18], or even chemical properties [19, 20, 21, 22, 23] non-intrinsically and on demand. The many-body physics originating from electronic and vibrational structure, as well as their respective decoherence, is crucial for a comprehensive understanding and refinement of this control-strategy.[24, 25, 26, 27, 28] Relevant experimental realizations include molecules or solids in micro- and plasmonic cavities, requiring a holistic treatment of electronic many-body physics coupled to various optical and phononic modes.[29, 16]
@@
Our primary challenge in this study is to address physical processes wherein many-body interactions, strong light-matter coupling, and strong system-environment coupling are all fundamental. Achieving this goal necessitates combining concepts from quantum optics and solid-state physics. Quantum optics often treats light-matter interaction non-perturbatively and offers many sophisticated techniques for dealing with strongly coupled environments [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], but it typically simplifies the structure of the materials drastically. Conversely, condensed matter or quantum chemistry approaches accurately describe the material system but oversimplify interactions with surrounding environments.[47, 11, 12] Here, we show how to bridge system-bath, quantum optical, and condensed matter approaches. We establish a framework for studying the effects of strongly coupled bosonic fields on many-body physics more efficiently than previously possible.
In order to deal with the strongly coupled environments we rely on the Hierarchical Equations Of Motion (HEOM) [45, 46] approach. HEOM has been used successfully to solve a large variety of problems [48, 49, 50, 51, 52, 53, 54, 55, 56]. It lends itself to our cause particularly well, because (i) its approximation of the environment is completely independent of the system dimension, and (ii) it can be modified to incorporate nonlinear terms (see Sec. II.2). However, it treats the system part exactly and thus fails for many-body systems due to the exponential scaling of the system dimension.
@@
Wavefunction-based approaches, such as tensor networks [57] or matrix product states, provide in principle exact and reliable results [58, 59, 60], but increase exponentially in cost over time. Their application is practically limited to short time scales of low-dimensional systems. Non-equilibrium Green functions approaches [61] made recent progress after the successful integration of the generalized Kadanoff-Baym ansatz [62, 63, 64] but a consistent treatment of complex structured baths remains challenging. An attractive approach with good scaling properties is to perform a cluster expansion of the density matrix into single-, two-, three-, and $N$-body reduced density matrices and then truncate the resulting Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy at some low order [65, 66].
Here, we combine BBGKY and HEOM to establish a stacked BBGKY-HEOM hierarchy that provides an efficient framework for the simulation of correlated many-body systems subject to structured baths. We apply our BBGKY-HEOM method to an experimentally relevant few-emitter lasing setup comprising multiple molecules inside a plasmonic cavity. There, we explicitly treat the vibrational modes of each molecule in a realistic manner by using parameters obtained from density-functional theory (DFT) calculations. Our results go beyond the state of the art and give new insights into how the vibrational spectrum influences the lasing process. Subsequently, we go past the paradigm of the two-level emitter and demonstrate that BBGKY-HEOM is able to capture many-body electronic dynamic in the Fermi-Hubbard model coupled to structured environments. BBGKY-HEOM provides accurate predictions as long as the onsite repulsion $U$, and with it the degree of electronic correlation, remains moderate. We notice that the bath adds a stabilizing contribution to the BBGKY equations.
@@
The remainder of this article is structured as follows: Sec. II provides a short introduction to non-Markovian quantum systems and many-body theory before the derivation of the BBGKY-HEOM equations in Sec. II.2. We illustrate their performance, limitation, and associated physical intuition for various systems in Sec. III. After benchmarking our method for many-emitter quantum optics in Sec. III.1, we showcase its potential by capturing vibrationally enhanced few-emitter lasing in Sec. III.2. We subsequently shift our focus to many-body electronic systems in Sec. III.3, where we investigate the Fermi-Hubbard model coupled to a lossy cavity mode and subject to charge noise . This system in then embedded into a highly structured bath in Sec. III.4 that represents phononic motion in organic crystals. Sec. IV finally concludes our study and provides an outlook into the near future.
@@ II Theory
A material comprising electrons and nuclei moving at non-relativistic velocities is described by Schrödinger’s equation. Its interplay with electromagnetic fields necessitates the consideration of the (quantized) normal modes of Maxwell’s equations. This combined system can be described using the minimal coupling Hamiltonian. In Coulomb gauge,
@@
for $N_{e}$ electrons and $N_{n}$ nuclei with charge $q_{i}$ and mass $m_{i}$. We will use atomic units in the following. Eq. (1) splits the transversal degrees from the longitudinal and instantaneous Coulombic interactions $H_{\parallel}$. The latter includes electronic and nuclear interactions which are initially treated in first quantization
@@
Their interplay results in rich many-body physics that gives rise to electronic, vibronic, and phononic structure and, unfortunately, to an exponentially increasing complexity. The open quantum systems approach now aims to identify a small subsystem of interest that we can treat explicitly (the ”system”) while absorbing the remainder into environments (the ”baths”) [67]. The first step is to simplify the Hamiltonian to a form that contains only linear interactions between system and bath(s)
@@
The gauge freedom of the electromagnetic modes provides various different paths for this. An intuitive approach is to isolate the bilinear term $\sum_{i}-i\hbar q_{i}/m_{i}\hat{\textbf{A}}(\textbf{r}_{i})\cdot\bm{\nabla}_{i}$ as coupling operator and absorb the quadratic diamagnetic term into dressed photonic operators [68, 69]. A second direction, often more promising for finite systems, is to shift into the multi-polar gauge (detailed in App. E). The combination with the long-wavelength approximation then provides a simple bilinear coupling between dipole moment and displacement field. In addition, self-polarization terms modify the dynamics of the system and the redefinition of the optical creation and annihilation operators introduce the need to reconsider the interpretation of system and bath. Vibrational and phononic interaction follow similarly. The influence of nuclear motion on the electron-nuclear interaction can be expanded around the equilibrium configuration
$\hat{V}_{en}\approx\hat{V}_{en}|_{0}+\sum_{j}\delta\textbf{R}_{j}\cdot\bm{\nabla}_{j}\hat{V}_{en}|_{0}+...$ (4)
and truncated at harmonic order, resulting in a bilinearly coupled electron-phonon or electron-vibration coupling that depends on the electronic configuration. Note that also here second-order (known as Debye-Waller) corrections appear that require to carefully scrutinize the separation between system and bath [70]. The following derivations and demonstrations are agnostic to such subtleties, as we will focus on the development and illustration of the BBGKY-HEOM methodology. A thorough derivation of system-bath couplings for a specific physical realization is evidently essential for meaningful predictions but it is not the focus of this manuscript.
@@ II.1 Many-body Open Quantum Systems beyond the Markovian Limit
Following the steps outlined above, both the phonon and the (transversal) photon sector of $\hat{H}_{mc}$ can be described as a collection of quantum harmonic oscillators that couple linearly to the electronic system. This description of an environment is frequently the case in quantum optics [71], or when the system is coupled weakly to a large number of degrees of freedom where central limit type of arguments yield an approximately Gaussian response [72]. We label these distinct environments with $k$, such that the bosonic annihilation (creation) operators belonging to mode $\lambda$ of environment $k$ are given by $a_{\lambda}^{k}$ ($a_{\lambda}^{k\dagger}$). The Hamiltonian corresponding to the total state $\rho_{tot}$ of system and environments is then
$H=$ $H_{sys}+\sum_{k,\lambda}g_{\lambda}^{k}(L^{k}a_{\lambda}^{k,\dagger}+L^{k\dagger}a^{k}_{\lambda})+\sum_{\lambda}\omega^{k}_{\lambda}a_{\lambda}^{k\dagger}a^{k}_{\lambda}.$ (5)
Here $L^{k}$ is the system operator that describes the coupling to the non-Markovian environment $k$ and $\omega_{\lambda}^{k}$, $g_{\lambda}^{k}$ are the frequencies and coupling strengths of the respective modes. The environments are characterized by their spectral density $J^{k}(\omega)=\sum_{\lambda}\lvert g^{k}_{\lambda}\rvert^{2}\delta(\omega-\omega^{k}_{\lambda})$ or equivalently by their bath correlation function $\alpha^{k}(\tau)=\int_{0}^{\infty}J_{k}(\omega)e^{-i\omega\tau}\mathrm{d}{\tau}$. When a description of the environment in terms of harmonic oscillators is suitable and the inital state of system and environment is a product state, the HEOM approach yields a numerically exact description for the evolution of the open quantum system, even for strong coupling [45, 46, 48]. To use HEOM, the bath correlation function needs to be fitted with exponential functions [73] according to
$\alpha^{k}(\tau)\approx\sum_{j}^{N_{exp}}G_{j}^{k}\exp\left(-w_{j}^{k}\tau\right),$ (6)
with $G_{j}^{k},\;w_{j}^{k}\in\mathbb{C}$. The resulting hierarchy couples the physical density matrix to auxiliary density matrices, and the number of auxiliary density matrices depends on the number of exponentials $N_{exp}$ as well as the system-environment coupling strength. In fact, HEOM type of techniques can be even extended to cases where the environment is anharmonic [74, 75]. In this work, we focus on the case of harmonic baths, either corresponding to optical modes confined in an imperfect cavity or a phonon bath into which a quantum emitter is embedded. The influence functional of the bath can then be computed exactly under the harmonic assumption by using the path integral formalism [76]. HEOM follows from this influence functional [46] and we provide a sketch of the derivation for the open system in App. B.
Following the previous motivation, we start with a many-body system that is linearly coupled to a collection of harmonic oscillators. The $N$-body system can be decomposed into a single particle Hamiltonian $H_{i}$ and two-particle interactions $V_{ij}$
$\begin{split}H_{sys}=&\sum_{i=1}^{N}H_{i}+\sum_{i\neq j}^{N}V_{ij}.\end{split}$ (7)
Additional (local) Markovian dissipative processes of the system are captured by $\mathcal{L}_{diss}$. The evolution of the total state $\rho_{tot}$ of system and environments is then described by
$\dot{\rho}_{tot}=$ $-i[H,\rho_{tot}]+\mathcal{L}_{diss}(\rho_{tot}).$ (8)
Eq. (8) together with Eq. (5)(7) provide the general form for systems discussed in this work. HEOM retains information about the bath, allowing one to access e.g. photonic correlation functions.
For the coupling operators $L_{k}$ we distinguish two cases, both illustrated in the following. First, all particles couple to the same environment with equal strength and the coupling operators take the form
$L^{k}=\sum_{i=1}^{N}L^{k}_{i}.$ (9)
@@
Second, each particle couples to its own local environment, however, all local environments are identical. This situation arises, for example, in a collection of molecules, each with its local vibrational bath (see Sec. III.2). The coupling operators are then denoted as $S_{k}$, and act only on the local Hilbert space of particle $k$. The evolution equation (5) is invariant under particle exchange. Such assumptions are natural for indistinguishable particles but one can relax this condition if one would like to describe ensembles of inhomogeneously broadened emitters.
Despite the restrictions employed above, the resulting system-bath Hamiltonian applies to a vast range of systems critical in fields such as quantum optics, quantum impurities in solid hosts, and strongly correlated materials in cavities.
@@ II.2 Combining HEOM with BBGKY
Let us now tackle the exponential growth of the Hilbert space by stacking HEOM and BBGKY hierarchy. We focus here on a single environment for the sake of simplicity, but note that the extension to multiple environments is trivial and will be illustrated in Sec. III.4. The HEOM equation accounts for the non-Markovianity of the coupled environment with an infinite set of coupled equations for the matrices $\rho^{(\bm{n},\bm{m})}$, whose dimension is equal to the dimension of the system Hilbert space. Here, $\rho^{(\bm{0},\bm{0})}=\operatorname{Tr}_{B}(\rho_{tot})=:\rho_{sys}$ is the physical density matrix of the system and $\bm{m},\bm{n}$ are vector indices that label auxiliary density matrices. The hierarchy is truncated at a sufficiently high order which is determined by the non-Markovianity of the system-bath interaction. A Markovian bath corresponds to a truncation at first order according to $\rho^{(\bm{e}_{k},\bm{0})}\propto L\rho^{(\bm{0},\bm{0})}$, $\rho^{(\bm{0},\bm{e}_{k})}\propto\rho^{(\bm{0},\bm{0})}L^{\dagger}$, where $\bm{e}_{k}$ denotes the unit vector in direction $k$. The evolution equation for the auxiliary density matrices (including the physical reduced density matrix $\rho^{(\bm{0},\bm{0})}$) for a single environment with coupling operator $L$ reads
$\dot{\rho}^{(\bm{n},\bm{m})}$ $=-i[H_{sys},\rho^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)\rho^{(\bm{n},\bm{m})}$
$+\sum_{j=1}^{N_{exp}}\bigg{(}G_{j}n_{j}L\rho^{(\bm{n}-\bm{e}_{j},\bm{m})}+G_{j}^{*}m_{j}\rho^{(\bm{n},\bm{m}-\bm{e}_{j})}L^{\dagger}$
$\quad\quad\quad+\left[\rho^{(\bm{n}+\bm{e}_{j},\bm{m})},L^{\dagger}\right]+\left[L,\rho^{(\bm{n},\bm{m}+\bm{e}_{j})}\right]\bigg{)}.$ (10)
Bold symbols indicate the vector character of a variable and the complex coefficients $G_{j},\bm{w}$ can be obtained from a fit according to Eq. (6) or inferred from scattering theory [77]. The application to many-body systems now requires us to identify a path to limit the dimensionality of the descriptor, so far $\rho^{(\bm{n},\bm{m})}$, to manageable levels.
@@
$\begin{split}F_{123}^{(\bm{n},\bm{m})}=&N(N-1)(N-2)\operatorname{Tr}_{4,5,...,N}(\rho^{(\bm{n},\bm{m})}),\\ F_{12}^{(\bm{n},\bm{m})}=&N(N-1)\operatorname{Tr}_{3,4,...,N}(\rho^{(\bm{n},\bm{m})}),\\ F_{1}^{(\bm{n},\bm{m})}=&N\operatorname{Tr}_{2,3,...,N}(\rho^{(\bm{n},\bm{m})}).\end{split}$ (11)
@@
Here $\operatorname{Tr}_{i,i+1,...,N}$ denotes the trace over all particles with index larger than $i$. The reduced density matrices are normalized to account for the particle number under contraction [78]. Note that since we are restricting ourselves to permutationally symmetric systems, the order of indexing is irrelevant.